# Improving the Euclidean Diffusion Generation of Manifold Data by Mitigating Score Function Singularity

### v1

@@ 8 Conclusion
The expansion (31) reduces to
$|\exp_{\tilde{x}}(v)-\tilde{x}|^{2}=|v|^{2}+O(|v|^{3})\,,$ (44)
and applying (28) we have
$(I-P(\exp_{\tilde{x}}(v)))(\exp_{\tilde{x}}(v)-\tilde{x})$ (45)
$=$ $\frac{1}{2}(I-P(\tilde{x}))\ddot{\gamma}_{\hat{v}}(0)|v|^{2}-\sum_{1\leq j,j^{\prime}\leq d}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3})$
$=$ $-\frac{1}{2}\sum_{1\leq j,j^{\prime}\leq n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3}),$
where we have used (26) and (27) to derive the first and the second equality, respectively. We denote $U\in\mathbb{R}^{n\times d}$ whose columns form an orthonormal basis of $T_{x}\mathcal{M}$. It is straightforward to verify that $U^{T}U=I\in\mathbb{R}^{d\times d}$ and $P(\tilde{x})=UU^{T}$. We can compute (43) as
@@
$=$ $(2\pi\sigma^{2})^{-\frac{n}{2}}\int_{\mathcal{O}}(p_{0}(\tilde{x})+O(|v|))\frac{1}{\sigma^{2}}\left(-\frac{1}{2}\sum_{1\leq j,j^{\prime}\leq n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3})\right)e^{-\frac{|v|^{2}+O(|v|^{3})}{2\sigma^{2}}}\mathrm{d}v$
$=$ $(2\pi\sigma^{2})^{\frac{d-n}{2}}\int_{\mathcal{O}_{\sigma}}(p_{0}(\tilde{x})+O(\sigma))\Big{[}-\frac{1}{2}\sum_{1\leq j,j^{\prime}\leq n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v^{\prime}_{j}v^{\prime}_{j^{\prime}}+O(\sigma)\Big{]}e^{-\frac{|v^{\prime}|^{2}+O(\sigma)}{2}}\mathrm{d}v^{\prime}$
$=$ $(2\pi\sigma^{2})^{\frac{d-n}{2}}\int_{\mathcal{W}_{\sigma}}\Big{[}-p_{0}(\tilde{x})\frac{1}{2}\sum_{1\leq j,j^{\prime}\leq n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})\Big{(}\sum_{l=1}^{d}U_{jl}w_{l}\Big{)}\Big{(}\sum_{l^{\prime}=1}^{d}U_{jl^{\prime}}w_{l^{\prime}}\Big{)}+O(\sigma)\Big{]}e^{-\frac{|Uw|^{2}+O(\sigma)}{2}}\mathrm{d}w$
$=$ $(2\pi\sigma^{2})^{\frac{d-n}{2}}\int_{\mathcal{W}_{\sigma}}\Big{[}-p_{0}(\tilde{x})\frac{1}{2}\sum_{1\leq j,j^{\prime}\leq n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})(UU^{T})_{jj^{\prime}}+O(\sigma)\Big{]}e^{-\frac{|w|^{2}+O(\sigma)}{2}}\mathrm{d}w$
$=$ $(2\pi\sigma^{2})^{\frac{d-n}{2}}\Big{[}-p_{0}(\tilde{x})\frac{1}{2}\sum_{j,j^{\prime}=1}^{n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma)\Big{]}\,,$
@@
$\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\frac{\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})}{p_{\sigma}(\tilde{x})}=\nabla^{\mathcal{M}}_{\tilde{x}}\log p_{0}(\tilde{x})-\frac{1}{2}\sum_{j,j^{\prime}=1}^{n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma).$ (47)
Finally, define $T(\tilde{x})=\sum_{j,j^{\prime}=1}^{n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})$, and we have
$T(\tilde{x})$ $=\sum_{j,j^{\prime}=1}^{n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}P_{jj^{\prime}}=\sum_{i,j,j^{\prime}=1}^{n}\frac{\partial(P_{\cdot i}P_{ij})}{\partial x_{j^{\prime}}}P_{jj^{\prime}}$ (48)
$=\sum_{i,j,j^{\prime}=1}^{n}\frac{\partial P_{\cdot i}}{\partial x_{j^{\prime}}}P_{ij}P_{jj^{\prime}}+P_{\cdot i}\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}P_{jj^{\prime}}=T(\tilde{x})+P(\tilde{x})T(\tilde{x}).$
This implies that $P(\tilde{x})T(\tilde{x})=0$. Therefore, multiplying $P(\tilde{x})$ on both sides of (47) confirms (12). Therefore, the second claim is obtained.
Appendix B Proof for Theorem 4.1
@@
We first study the properties of $\Sigma_{\sigma}(x)$. Recall that
@@
By the Sherman–Morrison–Woodbury Formula,
@@
$(x-\tilde{x})^{T}\Sigma_{\sigma}(x)^{-1}(x-\tilde{x})$ (52)
$=$ $\frac{1}{\sigma^{2}}(x-\tilde{x})^{T}P(x)(x-\tilde{x})+\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}(x-\tilde{x})^{T}(I-P(x))(x-\tilde{x})^{T}.$
We apply the same change of variables $x=\exp_{x^{*}}(v)$ as the proof in Section A and $x=x^{*}-\tilde{x}+v+O(|v|^{2})$. We obtain
$(x-\tilde{x})^{T}(I-P(x))(x-\tilde{x})^{T}$ (53)
@@
$(x^{*}-\tilde{x})^{T}(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k})(x^{*}-\tilde{x})$ (55)
$=$ $(x^{*}-\tilde{x})^{T}(\sum_{k}\frac{\partial(P)^{2}(x^{*})}{\partial x_{k}}v_{k})(x^{*}-\tilde{x})$
$=$ $(x^{*}-\tilde{x})^{T}P(x^{*})^{T}(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k})(x^{*}-\tilde{x})+(x^{*}-\tilde{x})^{T}(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k})P(x^{*})(x^{*}-\tilde{x})$
@@
$\tilde{S}_{jj^{\prime}}=\delta_{jj^{\prime}}+\sum_{i=1}^{n}(\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}})(x^{*})(x^{*}_{i}-\tilde{x}_{i})+\frac{1}{2}\sum_{k,l}\frac{\partial^{2}P_{kl}(x^{*})}{\partial x_{j}\partial x_{j^{\prime}}}(x^{*}_{k}-\tilde{x}_{l})(x^{*}_{k}-\tilde{x}_{l}),$ (56)
for $1\leq j,j^{\prime}\leq n$. By the Gershgorin circle theorem and the condition $\|x^{*}-\tilde{x}\|_{\infty}<\min\{1,\frac{2}{n^{2}(4M_{1}+M_{2})}\}$, the matrix $\tilde{S}$ is also positive definite. From (53)(54)(56), $p_{\sigma}(\tilde{x}|x)$ can be written as
@@
$\leq$ $\sigma^{m}\int_{\mathcal{M}}(2\pi)^{-\frac{n}{2}}|\det\Sigma_{\sigma}(x)|^{-\frac{1}{2}}|x-x^{*}|p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}-\frac{\delta}{\sigma^{2\alpha}}}$
@@
Similarly, the order of $-\Sigma_{\sigma}(x)^{-1}(\tilde{x}-x)$ can be estimated as
@@
$=$ $\frac{1}{\sigma^{2}}P(x)(x-\tilde{x})+\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}(I-P(x))(x-\tilde{x})$
$=$ $\frac{1}{\sigma^{2}}\big{(}P(x^{*})+\sum_{k}\frac{\partial P}{\partial x_{k}}(x^{*})v_{k}+O(|v|^{2})\big{)}\big{(}x^{*}-\tilde{x}+v+O(|v|^{2})\big{)}$
$+\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}(I-P(x^{*})-\sum_{k}\frac{\partial P}{\partial x_{k}}(x^{*})v_{k}+O(|v|^{2}))(x^{*}-\tilde{x}+v+O(|v|^{2}))$
@@
$\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})-\frac{x^{*}-\tilde{x}}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}p_{\sigma}(\tilde{x})$ (60)
@@
$=$ $(\sigma^{-\alpha})^{n-d}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}O(\sigma^{(1-2\alpha)\wedge 0}),$
@@
$p_{\sigma}(\tilde{x})=(\sigma^{-\alpha})^{n-d}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}\left(p_{0}(x^{*})+O(\sigma)\right).$ (61)
@@
$=$ $(2\pi\sigma^{2\alpha})^{\frac{d-n}{2}}p_{0}(\tilde{x})\Big{[}-\frac{1}{2}\sum_{j,j^{\prime}=1}^{n}\nabla_{\tilde{x}}P_{jj^{\prime}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\Big{]}.$
@@
$\quad\quad\cdot p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x).$
From (50), we have
[…]
Next, we continue to calculate (63) using (64) (65). There exists $\delta>0$ and $\mathcal{O}^{\prime}$ such that
@@
$=$ $(2\pi\sigma^{2\alpha})^{\frac{d-n}{2}}p_{0}(\tilde{x})\Big{[}-\frac{1}{2}\sum_{k,j,j^{\prime}}P_{\cdot k}\frac{\partial P_{jj^{\prime}}}{\partial x_{k}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\Big{]}+o(e^{-\frac{\delta}{4\sigma^{\alpha}}}).$
@@
$\sum_{j,j^{\prime}=1}^{n}\nabla P_{jj^{\prime}}P_{jj^{\prime}}=\sum_{k,j,j^{\prime}=1}^{n}\nabla(P_{jk}P_{kj^{\prime}})P_{jj^{\prime}}=$ $\sum_{k,j,j^{\prime}=1}^{n}\nabla(P_{kj^{\prime}})P_{jk}P_{jj^{\prime}}+\sum_{k,j,j^{\prime}=1}^{n}\nabla(P_{jk})P_{kj^{\prime}}P_{jj^{\prime}}$ (68)
$=$ $2\sum_{j,j^{\prime}=1}^{n}\nabla P_{jj^{\prime}}P_{jj^{\prime}}.$
implies $\sum_{j,j^{\prime}=1}^{n}\nabla P_{jj^{\prime}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})=0$. Therefore, (62) becomes
$\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})-\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi\sigma^{2\alpha})^{\frac{d-n}{2}}p_{0}(\tilde{x})O(\sigma^{1\wedge(2-2\alpha)}).$ (69)
@@
$\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi\sigma^{2\alpha})^{\frac{d-n}{2}}\left(\nabla_{\tilde{x}}^{\mathcal{M}}p_{0}(\tilde{x})+O(\sigma^{(2-2\alpha)\wedge 1})\right),$ (70)
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi\sigma^{2\alpha})^{\frac{d-n}{2}}\left(p_{0}(\tilde{x})+O(\sigma^{(2-2\alpha)\wedge 1})\right).$ (71)
Using (69) (70) (71), the second claim is obtained via

### v3

@@ 80. Guidelines:
In this case, we have $\tilde{x}\in\mathcal{M}$ and $\tilde{x}=x^{*}$. Therefore, the expansion (31) reduces to $|\exp_{\tilde{x}}(v)-\tilde{x}|^{2}=|v|^{2}+O(|v|^{3})$. Applying (28), we then obtain
[…]
$=$ $\frac{1}{2}(I-P(\tilde{x}))\ddot{\gamma}_{\hat{v}}(0)|v|^{2}-\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3})$
$=$ $-\frac{1}{2}\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3})\,$,
where we have used (26) and (27) to derive the first and the second equality, respectively. We denote $U\in\mathbb{R}^{n\times d}$ the matrix whose columns form an orthonormal basis of $T_{\tilde{x}}\mathcal{M}$. It is straightforward to verify that $U^{T}U=I\in\mathbb{R}^{d\times d}$ and $P(\tilde{x})=UU^{T}$. We can compute (43) as
@@
$=$ $(2\pi\sigma^{2})^{-\frac{n}{2}}\int_{\mathcal{O}}(p_{0}(\tilde{x})+O(|v|))\frac{1}{\sigma^{2}}\Big(-\frac{1}{2}\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3})\Big)e^{-\frac{|v|^{2}+O(|v|^{3})}{2\sigma^{2}}}\mathrm{d}v$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{d-n}\int_{\mathcal{O}_{\sigma}}(p_{0}(\tilde{x})+O(\sigma))\Big[-\frac{1}{2}\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v^{\prime}_{j}v^{\prime}_{j^{\prime}}+O(\sigma)\Big]e^{-\frac{|v^{\prime}|^{2}+O(\sigma)}{2}}\mathrm{d}v^{\prime}$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{d-n}\int_{\mathcal{W}_{\sigma}}\Big[-\frac{1}{2}p_{0}(\tilde{x})\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})\Big(\sum_{l=1}^{d}U_{jl}w_{l}\Big)\Big(\sum_{l^{\prime}=1}^{d}U_{j^{\prime}l^{\prime}}w_{l^{\prime}}\Big)$
$\hskip 256.0748pt+O(\sigma)\Big]e^{-\frac{|Uw|^{2}+O(\sigma)}{2}}\mathrm{d}w$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{d-n}\int_{\mathcal{W}_{\sigma}}\Big[-\frac{1}{2}p_{0}(\tilde{x})\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})(UU^{T})_{jj^{\prime}}+O(\sigma)\Big]e^{-\frac{|w|^{2}+O(\sigma)}{2}}\mathrm{d}w$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{d-n}\Big[-\frac{1}{2}p_{0}(\tilde{x})\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma)\Big]\,$,
@@
$\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\frac{\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})}{p_{\sigma}(\tilde{x})}=\nabla^{\mathcal{M}}_{\tilde{x}}\log p_{0}(\tilde{x})-\frac{1}{2}\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma).$ (46)
[…]
This implies that $P(\tilde{x})T(\tilde{x})=0$. Therefore, multiplying P(x~)P($\tilde${x}) on both sides of (46) confirms (12). The second claim is obtained.
Appendix B Proof of Theorem 4.1
@@
We first study the properties of $\Sigma_{\sigma}(x)$. Recall that for $x\in\mathcal{M}$,
@@
By the Sherman–Morrison–Woodbury formula [31],
@@
[…]
We apply the same change of variables $x=\exp_{x^{*}}(v)$ as in the proof in Section A and $x=x^{*}+v+O(|v|^{2})$. We obtain
[…]
@@
($x~)\displaystyle(x^{*}-\tilde{x})^{T}\left(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k}\right$)($^{*}-\tilde{x}$) (54)
$=$ ($x~)\displaystyle(x^{*}-\tilde{x})^{T}\left(\sum_{k}\frac{\partial(P)^{2}(x^{*})}{\partial x_{k}}v_{k}\right$)($^{*}-\tilde{x}$)
$=$ ($x~)\displaystyle(x^{*}-\tilde{x})^{T}P(x^{*})^{T}\left(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k}\right)(x^{*}-\tilde{x})+(x^{*}-\tilde{x})^{T}\left(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k}\right)P(x^{*}$)($^{*}-\tilde{x}$)
@@
$\tilde{S}_{jj^{\prime}}=\delta_{jj^{\prime}}+\sum_{i}\left(\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}\right)(x^{*})(x^{*}_{i}-\tilde{x}_{i})+\frac{1}{2}\sum_{k,l}\frac{\partial^{2}P_{kl}(x^{*})}{\partial x_{j}\partial x_{j^{\prime}}}(x^{*}_{k}-\tilde{x}_{l})(x^{*}_{k}-\tilde{x}_{l}),$ (55)
for $1\leq j,j^{\prime}\leq n$. By the Gershgorin circle theorem [41] and the condition $\|x^{*}-\tilde{x}\|_{\infty}<\min\{1,\frac{2}{n^{2}(4M_{1}+M_{2})}\}$, the matrix $\tilde{S}$ is also positive definite. From (52), (53), and (55), $p_{\sigma}(\tilde{x}|x)$ can be written as
@@
$\leq$ $\sigma^{m}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}-\frac{\delta}{\sigma^{2\alpha}}}\int_{\mathcal{M}}(2\pi)^{-\frac{n}{2}}|\det\Sigma_{\sigma}(x)|^{-\frac{1}{2}}|x-x^{*}|p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$
@@
Similarly, the order of $\Sigma_{\sigma}(x)^{-1}(\tilde{x}-x)$ can be estimated as
@@
[…]
$=$ $\frac{1}{\sigma^{2}}\left(P(x^{*})+\sum_{k}\frac{\partial P}{\partial x_{k}}(x^{*})v_{k}+O(|v|^{2})\right)\big(x^{*}-\tilde{x}+v+O(|v|^{2})\big)$
$+\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}\left(I-P(x^{*})-\sum_{k}\frac{\partial P}{\partial x_{k}}(x^{*})v_{k}+O(|v|^{2})\right)(x^{*}-\tilde{x}+v+O(|v|^{2}))$
@@
[…]
@@
$=$ $\sigma^{(d-n)\alpha}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}O(\sigma^{(1-2\alpha)\wedge 0}),$
@@
$p_{\sigma}(\tilde{x})=\sigma^{(d-n)\alpha}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}\left(p_{0}(x^{*})+O(\sigma)\right).$ (60)
@@
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}p_{0}(\tilde{x})\Big[-\frac{1}{2}\sum_{j,j^{\prime}=1}^{n}\nabla_{\tilde{x}}P_{jj^{\prime}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\Big].$
@@
$\quad\quad\cdot p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x),$
where $\Sigma_{\sigma}^{ij}$ refers to the $(i,j)$ entry of the matrix $\Sigma_{\sigma}$. From (49), we have
$(I-P(x))\Sigma_{\sigma}(x)^{-1}(x-\tilde{x})=\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}(I-P(x))(x-\tilde{x})=\frac{1}{\sigma^{2\alpha}}O(|v|^{2})$ (63)
and
$\frac{1}{2}\sum_{i,j,k}(x_{i}-\tilde{x}_{i})P_{\cdot k}\frac{\partial}{\partial x_{k}}(\Sigma_{\sigma}^{ij})^{-1}(x)(x_{j}-\tilde{x}_{j})$ (64)
$=$ $\frac{1}{2\sigma^{2}}\sum_{i,j,k}(x_{i}-\tilde{x}_{i})P_{\cdot k}\frac{\partial}{\partial x_{k}}P_{ij}(x)(x_{j}-\tilde{x}_{j})+\frac{1}{\sigma^{2\alpha}}O(|v|^{2})$
$=$ $\frac{1}{2\sigma^{2}}\sum_{i,j,k}P_{\cdot k}\frac{\partial P_{ij}}{\partial x_{k}}(\tilde{x})v_{i}v_{j}+\frac{1}{2\sigma^{2}}O(|v|^{3})+\frac{1}{\sigma^{2\alpha}}O(|v|^{2}).$
Next, we continue to calculate (62) using (63) and (64). There exists $\delta>0$ and $\mathcal{O}^{\prime}$ such that
@@
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}p_{0}(\tilde{x})\Big[-\frac{1}{2}\sum_{k,j,j^{\prime}}P_{\cdot k}\frac{\partial P_{jj^{\prime}}}{\partial x_{k}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\Big]+o(e^{-\frac{\delta}{4\sigma^{\alpha}}}).$
@@
$\sum_{j,j^{\prime}}\nabla P_{jj^{\prime}}P_{jj^{\prime}}=$ $\sum_{k,j,j^{\prime}}\nabla(P_{jk}P_{kj^{\prime}})P_{jj^{\prime}}$ (67)
$=$ $\sum_{k,j,j^{\prime}}\nabla P_{kj^{\prime}}P_{jk}P_{jj^{\prime}}+\sum_{k,j,j^{\prime}}\nabla P_{jk}P_{kj^{\prime}}P_{jj^{\prime}}$
$=$ $2\sum_{j,j^{\prime}}\nabla P_{jj^{\prime}}P_{jj^{\prime}}$
implies $\sum_{j,j^{\prime}}\nabla P_{jj^{\prime}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})=0$. Therefore, (61) becomes
$\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})-\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}p_{0}(\tilde{x})O(\sigma^{1\wedge(2-2\alpha)})\,.$ (68)
@@
$\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}\left(\nabla_{\tilde{x}}^{\mathcal{M}}p_{0}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\right),$ (69)
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}\left(p_{0}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\right).$ (70)
Using (68), (69), and (70), the second claim is obtained via